OpenStudy (anonymous):

Us the rule of 70 to estimate the years to double for an annual inflation rate of 7%

OpenStudy (dumbcow):

i thought it was the rule of 72

OpenStudy (dumbcow):

anyway you divide 70 by the interest rate ---> 70/7 = 10 10 yrs

OpenStudy (wolf1728):

I've heard it as the rule of 72 also.

OpenStudy (wolf1728):

By the way the exact figure for 7% to double in value is 10.2448 years.

OpenStudy (wolf1728):

70/7 = 10 72/7 = 10.2857142857 Seems the rule of 72 is more exact than the rule of 70 but the rule of 72 is NOT exact either.